please multiply this equation

(a+3)(2a-7)

thank you.

Answer by RAY100(1637)

(a+3)(2a-7)
,
2a^2 - 7a +6a -21
,
2a^2 - a - 21
check
let x=1, arbitrary
original,,,,(1+3)(2-7) = 4*(-5) = -20
answer,,, 2*1^2 -1 -21 = -20,,,,,ok